$SPAD/src/input parscurv.input

The Axiom Team December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1991.

———-

— * —

\getchunk{license}
-- test for ParametricSpaceCurve domain
draw(curve(tan(cos(t)),sin(t),cos(t)),t=-12..12)
draw(curve(sin(t),cos(t),sin(t)*cos(t)),t=-12..12)

———-

References^{[1] nothing}

4